# Formal extension of noncommutative tensor-triangular support varieties

### v1

@@ 1. Introduction
The first extended support varieties, designed to study infinitely-generated representations of a finite group over a field of positive characteristic, were introduced by Benson–Carlson–Rickard in [BCR95, BCR96], which used the Rickard idempotent functors developed in [Ric97]. This extended support variety theory was an extension of the rank / cohomological support. This approach was generalized by Benson–Iyengar–Krause to any triangulated category with an action by a graded-commutative Noetherian ring, which plays the role of the cohomology ring [BIK08, BIK12]. Shortly after the introduction of the Balmer support for symmetric tensor-triangulated categories, an analogous extension, also using versions of Rickard idempotents, was produced by Balmer–Favi [BF11]. This version of support was further developed and extended to tensor-triangulated module categories by Stevenson [Ste13, Ste18]. Beyond its use in classifying thick ideals, extended supports have also been used in the wide-ranging program of stratification, i.e.  in classifying localizing subcategories [BIK11, BIK11, Ste13, Ste14, Ste17, BIKP17, BIKP18, BHS23a, BHS23b].
@@ 2. Monoidal triangular geometry and support theories
Let $\mathbf{K}$ be a triangulated category, with shift functor $\Sigma$. For background on triangulated categories, see [Nee01]. Recall that a triangulated subcategory is one closed under cones and $\Sigma^{\pm}$; a thick subcategory is a full triangulated subcategory closed under direct summands; a localizing subcategory is a full triangulated subcategory closed under arbitrary set-indexed coproducts (if such coproducts exist).
@@ 3. Rickard idempotents and formal extensions of support
Let $\mathbf{K}$ be a big monoidal triangulated category, and let $\mathbf{I}$ be a thick ideal of $\mathbf{K}^{c}$. The construction of the Rickard idempotent functors ${\it{\Gamma}}_{\mathbf{I}}$ and $L_{\mathbf{I}}$ associated to $\mathbf{I}$ involves homotopy colimits; we recall the construction now. Suppose that
$A_{1}\xrightarrow{f_{1}}A_{2}\xrightarrow{f_{2}}A_{3}\xrightarrow{f_{3}}...$
is a sequence of maps between objects in $\mathbf{K}$. Denote by
$f:\bigoplus_{i=1}^{\infty}A_{i}\to\bigoplus_{i=1}^{\infty}A_{i}$
the map induced by the maps $f_{i}$. Then the homotopy colimit $\operatorname{hocolim}A_{i}$ of the sequence $\{f_{i}\}$ is defined as the cone of the map $1-f$, that is, there is a distinguished triangle
$\bigoplus_{i=1}^{\infty}A_{i}\xrightarrow{1-f}\bigoplus_{i=1}^{\infty}A_{i}\to\operatorname{hocolim}A_{i}.$
Let $A$ be some object of $\mathbf{K}$; denote $A_{0}:=A$. We denote by $f_{0}$ the natural map
$\bigoplus_{B\in\mathbf{I}}(B\otimes\operatorname{Hom}_{\mathbf{K}}(B,A_{0}))\to A_{0}.$
We now define inductively $A_{i+1}$ as the cone of $f_{i}$, where $f_{i}$ is defined analogously to $f_{0}$ but with $A_{i}$ replacing $A_{0}$. Since by construction we have a map $A_{i}\to A_{i+1}$, we get a map $g_{i}:A\to A_{i}$ via the composition
[…]
Completing the maps $g_{i}$ to distinguished triangles, we obtain
$B_{i}\to A\xrightarrow{g_{i}}A_{i}\to\Sigma B_{i}.$
There is a sequence of maps
$B_{1}\to B_{2}\to B_{3}\to...$
induced by the sequence
$A_{1}\to A_{2}\to A_{3}\to....$
We now set
${\it{\Gamma}}_{\mathbf{I}}A:=\operatorname{hocolim}B_{i},\;\;\;L_{\mathbf{I}}A:=\operatorname{hocolim}A_{i}.$
This process is functorial, and ${\it{\Gamma}}_{\mathbf{I}}$ and $L_{\mathbf{I}}$ can be described by the following universal property.
Theorem 3.1 ([Ric97]).
@@
with ${\it{\Gamma}}_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})$ and $L_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})^{\perp},$ in other words, there are no nonzero maps in $\mathbf{K}$ from any object of $\operatorname{Loc}(\mathbf{I})$ to $L_{\mathbf{I}}A.$
@@ 7. Faithfulness of extended support induced by a surjective map
Note that any comparative support is induced by a surjective map $\rho$; the main difference between the assumptions in this section and those previously in Section 5 is that we have dropped the assumption that $\sigma$ is tensorial, but have added the assumption that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian.
@@
It is clear that a support induced by a surjective map need not be tensorial. For instance, if $\operatorname{Spc}\mathbf{K}^{c}$ is the topological space consisting of two distinct closed points, $X$ is the one-point space, and $\rho$ is the continuous map $\operatorname{Spc}\mathbf{K}^{c}\to X$, then the support induced by $\rho$ is not tensorial.

### v2

@@ 1. Introduction
The first extended support varieties, designed to study infinitely-generated representations of a finite group over a field of positive characteristic, were introduced by Benson–Carlson–Rickard in [BCR95, BCR96], which used the Rickard idempotent functors developed in [Ric97]. This extended support variety theory was an extension of the rank / cohomological support. This approach was generalized by Benson–Iyengar–Krause to any triangulated category with an action by a graded-commutative Noetherian ring, which plays the role of the cohomology ring [BIK08, BIK12]. Shortly after the introduction of the Balmer support for symmetric tensor-triangulated categories, an analogous extension, also using versions of Rickard idempotents, was produced by Balmer–Favi [BF11]. This version of support was further developed and extended to tensor-triangulated module categories by Stevenson [Ste13, Ste18]. Beyond its use in classifying thick ideals, extended supports have also been used in the wide-ranging program of stratification, i.e.  in classifying localizing subcategories [BIK11, BIK11, Ste13, Ste14b, Ste17, BIKP17, BIKP18, BHS23a, BHS23b].
@@ 2. Monoidal triangular geometry and support theories
Let $\mathbf{K}$ be a triangulated category, with shift functor $\Sigma$. For background on triangulated categories, see [Nee01]. Recall that a triangulated subcategory is one closed under cones and $\Sigma^{\pm}$; a thick subcategory is a full triangulated subcategory closed under direct summands; a localizing subcategory is a full thick subcategory closed under arbitrary set-indexed coproducts (if such coproducts exist). We briefly remark that thickness follows automatically from the other properties of localizing subcategories as long as $\mathbf{K}$ is closed under set-indexed coproducts via the Eilenberg swindle, but the thickness hypothesis is important if such coproducts don’t exist.
@@ 3. Rickard idempotents and formal extensions of support
Let $\mathbf{K}$ be a big monoidal triangulated category, and let $\mathbf{I}$ be a thick ideal of $\mathbf{K}^{c}$. Then the Verdier quotient functor
$\mathbf{K}\to\mathbf{K}/\operatorname{Loc}(\mathbf{I})$
admits a right adjoint, by a general argument of Neeman (see [Nee92b, Lemma 1.7]) which was inspired by older arguments of Adams and Bousfield. By composing the quotient functor with its adjoint, one obtains a functor $L_{\mathbf{I}}:\mathbf{K}\to\mathbf{K}$ which takes value in $\operatorname{Loc}(\mathbf{I})^{\perp}$, the subcategory of $\mathbf{K}$ of objects $B$ such that $\operatorname{Hom}_{\mathbf{K}}(A,B)=0$ for all $A\in\operatorname{Loc}(\mathbf{I})$. By analogous argument, there exists a right adjoint to the inclusion functor
$\operatorname{Loc}(\mathbf{I})\to\mathbf{K},$
and again by composition we get a functor $\it{\Gamma}_{\mathbf{I}}:\mathbf{K}\to\mathbf{K}$ which takes value in $\operatorname{Loc}(\mathbf{I})$; see also [Kra10, §7.2]. A similar construction was given by Miller in [Mil92] for the category of spectra. Given an object $A$, the object $L_{\mathbf{I}}A$ can be described directly by a process of repeated homotopy colimits given by ranging over all possible maps from objects of $\mathbf{I}$. This construction was also used directly by Rickard in [Ric97], who subsequently used the functors $\it{\Gamma}_{\mathbf{I}}$ and $L_{\mathbf{I}}$ in joint work with Benson and Carlson to define the first extended supports in modular representation theory [BCR95, BCR96]. The functors $\it{\Gamma}_{\mathbf{I}}$ and $L_{\mathbf{I}}$, which are now commonly referred to as Rickard idempotent functors, satisfy the following universal property:
Theorem 3.1 ([Nee92b, Ric97]).
@@
with ${\it{\Gamma}}_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})$ and $L_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})^{\perp}.$
@@
The results in this paper will focus on support data based on Noetherian spaces $X$. For non-Noetherian spaces, one can define $\it{\Gamma}_{x}$ in a more general way using Thomason subsets, as discussed in [Ste14a, §2.1], to get around the issue that $\mathcal{V}(x)$ might not be Thomason. Recall that in a Noetherian space, any specialization closed subset is automatically a Thomason subset.
Remark 3.5.
@@ 7. Faithfulness of extended support induced by a surjective map
Note that any comparative support is induced by a surjective map, since it is induced by the comparison map $\rho$ by Lemma 5.2, parts (1) and (2); the main difference between the assumptions in this section and those previously in Section 5 is that we have dropped the assumption that $\sigma$ is tensorial, but have added the assumption that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian. The purpose is for eventual application to the central cohomological support in the subsequent section (Section 8); for a finite tensor category $\mathbf{C}$, both the questions of Noetherianity of $\operatorname{Spc}\underline{\mathbf{C}}$ and the tensoriality of the central support are open.
@@
It is clear that a support induced by a surjective map need not be tensorial. For instance, if $\operatorname{Spc}\mathbf{K}^{c}$ is the topological space consisting of two distinct closed points, $X$ is the one-point space, and $\rho$ is the unique continuous map $\operatorname{Spc}\mathbf{K}^{c}\to X$. Then the support induced by $\rho$ is not tensorial: if $x$ is the unique point of $X$, and if $A$ and $C$ are two objects of $\mathbf{K}^{c}$ with disjoint nonempty supports (which exist, by Theorem 2.9), then we have $\sigma(A)=\{x\}=\sigma(C)$, but also have
$\bigcup_{B\in\mathbf{K}^{c}}\operatorname{supp}^{\rm B}(A\otimes B\otimes C)=\operatorname{supp}^{\rm B}(A)\cap\operatorname{supp}^{\rm B}(C)=\varnothing,$
hence $A\otimes B\otimes C\cong 0$ for all objects $B$. But then
$\bigcup_{B\in\mathbf{K}^{c}}\sigma(A\otimes B\otimes C)=\varnothing\subsetneq\sigma(A)\cap\sigma(B)=\{x\}.$